OpenStudy (anonymous):

Use graphs and tables to find the limit and identify any vertical asymptotes of limit of 1 divided by the quantity x minus 9 as x approaches 9 from the left

OpenStudy (mathmate):

 The above shows the graph y=1/x for f(x)=1/(x-9), it's the same graph translated to the right by 9 units.

OpenStudy (anonymous):

so its c?

OpenStudy (anonymous):

@mathmate ?

OpenStudy (mathmate):

"for f(x)=1/(x-9), it's the same graph translated to the \(right\) by 9 units."

OpenStudy (anonymous):

so than -infinite; x=9

OpenStudy (mathmate):

Yes, as explained below.  When f(x) is transformed to g(x)=a.f(x-h)+k, it is vertically stretched by "a", translated to the right by h, and translated upwards by k. Here, a=1, k=0 and h=9, so it moves to the right by 9 units.